Normal saline solution used in IVs is usually around 0.90% m/v or 0.154 M; however, other saline solutions may have a higher or lower concentration of sodium chloride, NaCl, depending on the use. For example, a hypertonic saline solution with a concentration greater than 0.154 M may be used to treat hyponatremia (sodium concentration in the serum is too low).

Calculate the molarity of the NaCl solution when 5.84 g of NaCl is dissolved in 100 mL of water.
Express the molarity with the appropriate units.
1
Expert's answer
2015-11-04T10:04:46-0500
By definition, the molarity of the solution is the number of the moles of the solute divided by the volume of the solution. The number of the moles of NaCl can be calculated from its mass and molar mass:
c = n/V = m/MV = 5.84g/(58.5g/mol * 0.1L) = 1 mol/L

Answer: 1mol/L